the length of a rectangular garden is 7 feet longer than its width. the gardens perimeter is 186. find the length of the garden.
Answer by algebrahouse.com(1659)

x = width
x + 7 = length {length is 7 more than width}

Perimeter of a rectangle = 2(width) + 2(length)

2x + 2(x + 7) = 186 {substituted width, length, and area into perimeter formula}
2x + 2x + 14 = 186 {used distributive property}
4x + 14 = 186 {combined like terms}
4x = 172 {subtracted 14 from each side}
x = 43 {divided each side by 4}
x + 7 = 50 {substituted 43, in for x, into x + 7}

length of garden is 50 ft

For more help from me, visit: www.algebrahouse.com
Also, you can Follow @AlgebraHouse on Twitter